Two blocks, each with weight w, are held in place on a frictionless incline as shown below. In
terms of w and the angle of the incline, calculate the tension in (a) the rope connecting the blocks and (b) the rope that connects block A to the wall. (c) Calculate the magnitude of the force that the incline exerts on each block. (d) Interpret your answers for the cases in which the angles are 0 degrees & 90 degrees
1
Expert's answer
2020-04-20T10:44:05-0400


a)


"T_1=m_Bg\sin{\alpha}=W\sin{\alpha}"

b)


"T_2=T_1+m_Ag\sin{\alpha}=W\sin{\alpha}+W\sin{\alpha}\\=2W\sin{\alpha}"

c)


"R_A=m_Ag\cos{\alpha}=W\cos{\alpha}"

"R_B=m_Bg\cos{\alpha}=W\cos{\alpha}"

d) 1) "\alpha=0\degree"

"T_1=T_2=0\ N"

"R_A=R_B=W"

We have two blocks on horizontal plane, no forces between them.

2) "\alpha=90\degree"


"R_A=R_B=0\ N"

"T_1=W, T_2=2W"

Both blocks are hanging in the air.